A university administrator would like to estimate the true mean salary of all professors at the university. She uses $4000 as the population standard deviation of the salary of all professors at the university. What sample size would be required so that a 90% confidence interval for 𝜇 has a length of $2000?


1
Expert's answer
2021-02-25T00:45:37-0500

Given,

Standard deviation "\sigma=" $"4000"

Mean salary "\mu=" $"2000"

Confidence level="90%" %

Let "n" be the sample size

The value of z at 90% confidence level is 1.645


As we know, "z=\dfrac{x-\mu}{\sigma}"


"1.645=\dfrac{x-2000}{4000}"


"x=1.64\times4000+2000=8560" $


Also The confidence interval is given by-

"x=\mu\pm z\dfrac{\sigma}{\sqrt{n}}"


"\implies 8560=2000\pm 1.645\times \dfrac{4000}{\sqrt{n}}"


"\implies \dfrac{4000}{\sqrt{n}}=\dfrac{8560-2000}{1.645}"


"\implies \sqrt{n}=\dfrac{4000\times 1.645}{6560}=1.003"


"\implies n=1.01"